# SR-Scientist: Scientific Equation Discovery With Agentic AI

### v1

@@ Acknowledgments
Appendix A Appendix
A.1 Baseline Implementation Details
@@
LLM-SR (Shojaee et al., 2024) leverages the extensive scientific knowledge and code-generation capabilities of LLMs to discover mathematical equations. It treats equations as programs and combines an LLM’s scientific priors with an evolutionary search to find accurate and interpretable formulas. We implement it using the llm-srbench [^10] https://github.com/deep-symbolic-mathematics/llm-srbench repository with its default configuration and set the maximum number of LLM calls to 1,000 for each problem.
A.2 Details of SR-Scientist Inference Framework
For each iteration, we set the initial MAPE goal to 0.1% and the termination MAPE threshold to 0.0001%. For each LLM, we set the sampling temperature to 0.7 and the maximum completion length per call to 8,192. For the equation evaluator tool, to control the length and the complexity of the generated equations, we set the maximum number of parameters to 10 in all experiments, following Shojaee et al. (2024). To ensure a fair comparison with methods like LLM-SR, the optimization goal for the BFGS algorithm is the mean squared error. For parameter optimization, we use the scipy library for nonlinear optimization with the BFGS algorithm. The prompt for the LLM agent is shown in Figure 13 and Figure 14. For Qwen3-Coder-480B and GLM-4.5-Air, we use their FP8 quantization versions.
A.3 Details of SR-Scientist Training Framework
A.3.1 Training Data Synthesis
We synthesize our training data following the approach of Shojaee et al. (2024), using a hybrid model-based and rule-based method. The problems cover four scientific disciplines: materials science (Stress with respect to Strain and Temperature), chemistry (Reaction rate with respect to Time and Concentration), biology (Growth rate with respect to Time and Population size), and physics (Acceleration with respect to Time, Displacement, and Velocity). The detailed procedure is as follows:
@@ A.3.1 Training Data Synthesis
    Data Point Generation: For the material science equations, which represent static systems, we generate data points by uniformly sampling across a defined range of parameters. The temperature $T$ is sampled from the range of $273\penalty 10000\ \text{K},573\penalty 10000\ \text{K}]$, and strain $\epsilon$ is sampled from $[0,0.6]$. A total of 5000 points are generated using an evenly spaced grid. The OOD test set is created by taking the 500 data points with the highest temperature values. The remaining 4500 points are then randomly shuffled, with 500 points allocated to the test set and the rest forming the training set. For the chemistry, biology, and physics equations, which represent dynamic systems, we use the solve_ivp function from the SciPy Python package with the RK45 method to solve the differential equations. Given initial conditions at $t=0$, we solve the equations to obtain the relationship between variables over time (equations that cannot be solved are simply discarded). We generate 5000 data points by uniformly sampling the time variable $t$ from the range of $[0,60]$. The OOD test set is created by selecting the 500 data points with the highest time values. The remaining 4500 points are randomly partitioned into a 500-point test set and the remaining training set.
@@ A.3.3 Training Details
A.3.4 The Effect of LLM Backbone
To further investigate the effect of RL training on different models, we conduct RL training on the Qwen3-8B, Qwen3-14B, and Qwen3-32B models using the same training configuration detailed in Appendix A.3.3 and present the results in Table 5. All models achieve significant performance gains after RL training, with most improving their performance across all subjects. When comparing LLM-SR and SR-Scientist without additional RL training, SR-Scientist outperforms LLM-SR on $Acc_{\text{0.01}}$ and matches it on $Acc_{\text{0.1}}$ when using Qwen3-32B as the LLM backbone. However, it lags behind when using the Qwen3-8B and Qwen3-14B models. This highlights the importance of a model’s intrinsic ability in a framework that provides the model with greater autonomy.
@@ A.3.4 The Effect of LLM Backbone
A.4 Analysis for the evaluation metrics
@@ A.5 Cost
A.6 OOD Performance of Other Models
We illustrate the OOD performance using other models as LLM backbones in Figures 9, 10, 11, and 12. For models including GLM-4.5-Air, GPT-OSS-120B, and GPT-OSS-20B, SR-Scientist consistently outperforms the other methods in overall accuracy and most subjects. For Qwen3-Coder-30B, it slightly lags behind the method PySR on the OOD data. After RL training, it not only enhances its performance on ID data but also on OOD data, showing the generalization of the discovered equations.
@@ A.6 OOD Performance of Other Models
A.7 Symbolic Accuracy Calculation
@@
A.8 Tool Call Analysis

### v2

@@ Acknowledgments
Appendix
Appendix A Baseline Implementation Details
@@
LLM-SR (Shojaee et al., 2025a) leverages the extensive scientific knowledge and code-generation capabilities of LLMs to discover mathematical equations. It treats equations as programs and combines an LLM’s scientific priors with an evolutionary search to find accurate and interpretable formulas. We implement it using the llm-srbench [^12] https://github.com/deep-symbolic-mathematics/llm-srbench repository with its default configuration and set the maximum number of LLM calls to 1,000 for each problem.
Appendix B Details of SR-Scientist Inference Framework
For each iteration, we set the initial MAPE goal to 0.1% and the termination MAPE threshold to 0.0001%. For each LLM, we set the sampling temperature to 0.7 and the maximum completion length per call to 8,192. For the equation evaluator tool, to control the length and the complexity of the generated equations, we set the maximum number of parameters to 10 in all experiments, following Shojaee et al. (2025a). To ensure a fair comparison with methods like LLM-SR, the optimization goal for the BFGS algorithm is the mean square error. For parameter optimization, we use the scipy library for nonlinear optimization with the BFGS algorithm. The prompt for the LLM agent is shown in Figure 15 and Figure 16.
Appendix C Details of SR-Scientist Training Framework
C.1 Training Data Synthesis
We synthesize our training data following the approach of Shojaee et al. (2025a), using a hybrid model-based and rule-based method. The problems cover four scientific disciplines: materials science (Stress with respect to Strain and Temperature), chemistry (Reaction rate with respect to Time and Concentration), biology (Growth rate with respect to Time and Population size), and physics (Acceleration with respect to Time, Displacement, and Velocity). The detailed procedure is as follows:
@@ C.1 Training Data Synthesis
    Data Point Generation: For the material science equations, which represent static systems, we generate data points by uniformly sampling across a defined range of parameters. The temperature $T$ is sampled from the range of $273~\text{K},573~\text{K}]$, and strain $\epsilon$ is sampled from $[0,0.6]$. A total of 5000 points are generated using an evenly spaced grid. The OOD test set is created by taking the 500 data points with the highest temperature values. The remaining 4500 points are then randomly shuffled, with 500 points allocated to the test set and the rest forming the training set. For the chemistry, biology, and physics equations, which represent dynamic systems, we use the solve_ivp function from the SciPy Python package with the RK45 method to solve the differential equations. Given initial conditions at $t=0$, we solve the equations to obtain the relationship between variables over time (equations that cannot be solved are simply discarded). We generate 5000 data points by uniformly sampling the time variable $t$ from the range of $[0,60]$. The OOD test set is created by selecting the 500 data points with the highest time values. The remaining 4500 points are randomly partitioned into a 500-point test set and the remaining training set.
@@ C.3 Training Details
C.4 The Effect of LLM Backbone
To further investigate the effect of RL training on different models, we conduct RL training on the Qwen3-8B, Qwen3-14B, and Qwen3-32B models using the same training configuration detailed in Appendix C.3 and present the results in Table 5. All models achieve significant performance gains after RL training, with most improving their performance across all subjects. When comparing LLM-SR and SR-Scientist without additional RL training, SR-Scientist outperforms LLM-SR on $Acc_{\text{0.01}}$ and matches it on $Acc_{\text{0.1}}$ when using Qwen3-32B as the LLM backbone. However, it lags behind when using the Qwen3-8B and Qwen3-14B models. This highlights the importance of a model’s intrinsic ability in a framework that provides the model with greater autonomy.
@@ C.4 The Effect of LLM Backbone
Appendix D Supplementary Analysis and Evaluation Details
D.1 Analysis for the evaluation metrics
@@ D.2 Cost
D.3 OOD Performance of Other Models
We illustrate the OOD performance using other models as LLM backbones in Figures 9, 12, 12, and 12. For models including GLM-4.5-Air, GPT-OSS-120B, and GPT-OSS-20B, SR-Scientist consistently outperforms the other methods in overall accuracy and most subjects. For Qwen3-Coder-30B, it slightly lags behind the method PySR on the OOD data. After RL training, it not only enhances its performance on ID data but also on OOD data, showing the generalization of the discovered equations.
@@ D.3 OOD Performance of Other Models
D.4 Other Qualitative Results
To demonstrate how the agent gains insight from data analysis and long-horizon exploration, we provide two snippets of the agent trajectories in Figure 13 and Figure 14. As shown by the reasoning process in the dashed box, the agent forms its hypothesis for the equation based on its analysis of the data through code, along with the performance of historical equations. This analysis grounded in the data helps decrease the search space compared to previous methods. The reasoning and analysis process can also inspire human scientists towards better equation design.
$K$ $\text{Acc}_{0.01}$ $\text{Acc}_{0.001}$
1 63.57 49.61
3 63.57 49.35
5 63.31 50.90
10 61.50 47.29
Table 9: Sensitivity analysis of the number of equations $K$ fetched from the experience buffer. (with GPT-OSS-120B as LLM backbone)
$s_{\text{goal}}$ $\text{Acc}_{0.01}$ $\text{Acc}_{0.001}$
0.1 64.34 48.84
0.01 63.31 49.61
0.001 63.57 49.35
0.0001 51.74 37.70
Table 10: Sensitivity analysis of the initial MAPE goal $s_{\text{goal}}$. (with GPT-OSS-120B as LLM backbone)
D.5 Sensitivity Analysis of Hyperparameters
We present a sensitivity analysis of the hyperparameters in Table 9 and Table 10, covering the number of equations fetched from the experience buffer and the initial MAPE goal. Regarding the number of fetched equations, our method remains robust when this value is relatively small. However, performance declines slightly when it is increased to 10. This suggests that including a large number of fetched equations in the context window may strain the model’s capability. Regarding the initial MAPE goal, performance is insensitive to relatively easy targets, as our dynamic goal update mechanism effectively prevents trivial goals. However, an extremely challenging goal (e.g., 0.0001) leads to degraded performance, causing the model to become trapped during the optimization process.
D.6 Symbolic Accuracy Calculation
@@
D.7 Tool Call Analysis
@@
Appendix E The Use of Large Language Models
The large language model was utilized as a writing assistant during the preparation of this manuscript. Its role was strictly limited to proofreading for grammatical errors, improving sentence structure, and enhancing readability. The large language model was not used for generating any core ideas, methodology, or scientific content. The authors take full responsibility for all content presented.
@@ Appendix E The Use of Large Language Models
Agent: Now I’$l$ analyze the data further to understand the relationship better. Let me look at how the growth rate changes with respect to population and time. ‘‘‘ import json import sys import numpy as np # load the data input_data_str = sys.stdin.read() data_list = json.loads(input_data_str) # Convert to numpy arrays data_array = np.array(data_list) outputs = data_array[:, 0] # Population growth rate t_values = data_array[:, 1] # Time P_values = data_array[:, 2] # Population # Look at the pattern when population is high (near carrying capacity) high_pop_mask = P_values > 90 high_pop_data = data_array[high_pop_mask] print(f"Number of high population (>90) data points: {len(high_pop_data)}") if len(high_pop_data) > 0: high_pop_growth_rates = high_pop_data[:, 0] print(f"High population growth rates: min={np.min(high_pop_growth_rates):.6f}, max={np.max(high_pop_growth_rates):.6f}, mean={np.mean(high_pop_growth_rates):.6f}") # Look at the pattern when population is low low_pop_mask = P_values < 20 low_pop_data = data_array[low_pop_mask] print(f"Number of low population (<20) data points: {len(low_pop_data)}") if len(low_pop_data) > 0: low_pop_growth_rates = low_pop_data[:, 0] print(f"Low population growth rates: min={np.min(low_pop_growth_rates):.6f}, max={np.max(low_pop_growth_rates):.6f}, mean={np.mean(low_pop_growth_rates):.6f}") # Look at early time vs late time patterns early_time_mask = t_values < 10 late_time_mask = t_values > 40 early_data = data_array[early_time_mask] late_data = data_array[late_time_mask] print(f"Early time growth rates: min={np.min(early_data[:, 0]):.6f}, max={np.max(early_data[:, 0]):.6f}, mean={np.mean(early_data[:, 0]):.6f}") print(f"Late time growth rates: min={np.min(late_data[:, 0]):.6f}, max={np.max(late_data[:, 0]):.6f}, mean={np.mean(late_data[:, 0]):.6f}") # Check if there’s a quadratic relationship P_squared = P_values ** 2 correlation_p2_o = np.corrcoef(P_squared, outputs)[0, 1] print(f"Correlation between P^2 and Growth Rate: {correlation_p2_o:.6f}") ‘‘‘ Tool: Number of high population (>90) data points: 3468 High population growth rates: min=-0.051324, max=5.084093, mean=0.118920 Number of low population (<20) data points: 234 Low population growth rates: min=1.016704, max=16.429510, mean=6.147035 Early time growth rates: min=0.792945, max=24.778742, mean=10.340174 Late time growth rates: min=-0.048202, max=0.077139, mean=0.008463 Correlation between P^2 and Growth Rate: -0.704877 ------------------------------------------------------------------------------------------------------ Agent: Based on my analysis, I can see that this appears to be a population growth model with characteristics of logistic growth: High population values (>90) have very low growth rates, approaching zero. Low population values (<20) have high growth rates. Early times have higher growth rates than late times. There’s a strong negative correlation between P^2 and growth rate. This suggests a logistic growth model where the growth rate decreases as population approaches a carrying capacity. Let me try a more refined logistic model: ------------------------------------------------------------------------------------------------------ def equation(t: np.ndarray, P: np.ndarray, params: np.ndarray) -> np.ndarray: r, K, alpha, beta = params[0], params[1], params[2], params[3] # Ensure parameters stay in reasonable ranges K = np.abs(K) + 90.0 # Near maximum observed population alpha = np.abs(alpha) # Ensure positive decay rate logistic_term = r * P * (1 - P / K) time_decay = np.exp(-alpha * t) output = logistic_term * time_decay + beta return output
Figure 13: The snippet of the agent trajectory for the BPG18 biology problem. (with Qwen3-Coder-480B as LLM backbone)
[…]
Figure 14: The snippet of the agent trajectory for the PO29 physics problem. (with GPT-OSS-120B as LLM backbone)